OpenStudy (anonymous):

can someone explain to me how to do this please? Let f(x) = x + 8 and g(x) = x^2 – 6x – 7. Find f(g(2)). –7 –3 10 33

OpenStudy (ajprincess):

First find f(g(x)). f(g(x))=(x^2 – 6x – 7.)+8 can u do?

OpenStudy (anonymous):

what do i do first?

OpenStudy (anonymous):

confused!

OpenStudy (ajprincess):

f(g(x))=(x^2 – 6x – 7.)+8 =x^2-6x-7+8 =x^2-6x+1 Nw f(g(2))=(2)^2-6(2)+1 can u do nw?

OpenStudy (anonymous):

i got -7

OpenStudy (ajprincess):

That's right Let f(x)=2x+1 and g(x)=x+1 To find f(g(x)) jst substitute g(x) in place of x in f(x). That is f(g(x))=2(x+1)+1 Nw r u k with finding f(g(x))?

OpenStudy (anonymous):

i got it but i just don't know how you set it up

OpenStudy (ajprincess):

vch one?

OpenStudy (anonymous):

(x^2 – 6x – 7.)+8

OpenStudy (ajprincess):

Did u understand my example?

OpenStudy (anonymous):

no

OpenStudy (ajprincess):

ok. Let us assume that f(x)=2x and g(x)=3x. k?

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

 Is it clear?